# ViLA: Efficient Video-Language Alignment for Video Question Answering

### v3

@@ Abstract
In this work, we propose an efficient Video-Language Alignment (ViLA) network. Our ViLA model addresses both efficient frame sampling and effective cross-modal alignment in a unified way. In our ViLA network, we design a new learnable text-guided Frame-Prompter together with a cross-modal distillation (QFormer-Distiller) module. Pre-trained large image-language models have shown promising results on problems such as visual question answering (VQA). However, how to efficiently and effectively sample video frames when adapting pre-trained large image-language model to video-language alignment is still the major challenge. Compared with prior work, our ViLA model demonstrates the capability of selecting key frames with critical contents, thus improving the video-language alignment accuracy while reducing the inference latency (+3.3% on NExT-QA Temporal with 3.0$\times$ speed up). Overall, our ViLA network outperforms the state-of-the-art methods on the video question-answering benchmarks: +4.6% on STAR Interaction, +2.2% on STAR average with 3.0$\times$ speed up, ours 2-frames out-perform SeViLA 4-frames on the VLEP dataset with 4.2$\times$ speed-up.
@@ 1 Introduction
Figure 2: Model Overview. Our ViLA model includes 4 sub-modules: the visual encoder, text-supervised Frame-Prompter (FP), QFormer-Distiller (QFD), and a LLM. We encode the video frames through a frozen visual encoder. Then we train the Teacher-QFormer using all the frame features. After that, we train the Student-QFormer and Frame-Prompter end-to-end. Unlike the Teacher-QFormer, our Student-QFormer is trained with masked frames features from a text-supervised Frame-Prompter. Finally, the input question text and QFormer transformed visual features go through a frozen large language model to generate the answer. Our network supports both leveraging LLM through proper visual prompting without affecting the original LLM (Frozen) ability on language tasks and fineuning LLMs(LoRA) simultaneously to get optimal performance on specific tasks.
@@ 2.1 Visual-Language Alignment
One major downstream task of Video-Language pre-training is Video Question Answering (VQA). Early works [2, 22, 63] often rely on human annotated datasets, while recent works [58, 67, 66] make better use of large-scale data from public. Along with the quick advances in VQA methods, A lot of benchmark datasets have also been introduced to standardize the model performance comparison, including NExT-QA [56], STAR [54], How2QA [31], TVQA [25], and VLEP [26]. We benchmark our network mainly on the Video Question and Answering task. It’s also the major task to test the effectiveness of the video-language alignment.
@@ 3.2 Text-guided Frame-Prompter Learning
$\pi=\left\{p_{k}\mid p_{k}=\frac{\exp(s_{k})}{\sum_{j=1}^{T/S}\exp(s_{j})},k,j\in[1,T/S]\right\},\vspace{-2mm}$ (3)
where $S$ is the number of frames of a segment. Then we compute our segment mask $M_{S}$ using the segment probability $p_{k}$ and the $g_{k}$ is sampled from the $\text{Gumbel}(0,1)$ distribution,
$M_{S}=\mathcal{H}\left(\arg\max_{k}\left[g_{k}+\log p_{k}\right]\right)k\in[1,T/S],\vspace{-2mm}$ (4)
@@ 5 Conclusion, Limitation and Future Work
How to properly ingest visual information to LLMs to utilize its capability effectively in cross-modal tasks? In this work, we present a new ViLA network with a new text-guided Frame-Prompter to smartly sample important frames, together with a cross-modal temporal distillation (QFormer-Distiller) for efficient and effective video-language alignment. From our experiments, our ViLA out-performs SOTA methods on four video question answering benchmarks and one video event prediction benchmark, especially on the temporal (including Temporal, Causal, Interaction and etc.) type of questions. For the limitation, we don’t have a change to evaluate on LLMs larger than 13b due to the resource constraints. We plan to continue research on the design of our Frame-Prompter, especially on long video-language alignment.

### v4

@@ Abstract
We propose an efficient Video-Language Alignment (ViLA) network. Our ViLA model addresses both efficient frame sampling and effective cross-modal alignment in a unified way. In our ViLA network, we design a new learnable text-guided Frame-Prompter together with a cross-modal distillation (QFormer-Distiller) module. Pre-trained large image-language models have shown promising results on problems such as visual question answering (VQA). However, how to efficiently and effectively sample video frames when adapting pre-trained large image-language model to video-language alignment is still the major challenge. Compared with prior work, our ViLA model demonstrates the capability of selecting key frames with critical contents, thus improving the video-language alignment accuracy while reducing the inference latency (+3.3% on NExT-QA Temporal with 3.0$\times$ speed up). Overall, our ViLA network outperforms the state-of-the-art methods on the video question-answering benchmarks: +4.6% on STAR Interaction, +2.2% on STAR average with 3.0$\times$ speed up, ours 2-frames out-perform SeViLA 4-frames on the VLEP dataset with 4.2$\times$ speed-up. Code will be available at https://github.com/xijun-cs/ViLA.
@@ 1 Introduction
Figure 2: Model Overview. Our ViLA model includes 4 sub-modules: the visual encoder, text-supervised Frame-Prompter (FP), QFormer-Distiller (QFD), and a LLM. We encode the video frames through a frozen visual encoder. Then we train the Teacher-QFormer using all the frame features. After that, we train the Student-QFormer and Frame-Prompter end-to-end. Unlike the Teacher-QFormer, our Student-QFormer is trained with masked frames features from a text-supervised Frame-Prompter. Finally, the input question text and QFormer transformed visual features go through a frozen large language model to generate the answer. Our network supports both leveraging LLM through proper visual prompting without affecting the original LLM (Frozen) ability on language tasks and finetuning LLMs(LoRA) simultaneously to get optimal performance on specific tasks.
@@ 2.1 Visual-Language Alignment
One major downstream task of Video-Language pre-training is Video Question Answering (VQA). Early works [2, 22, 63] often rely on human annotated datasets, while recent works [58, 67, 66] make better use of large-scale data from public. Along with the quick advances in VQA methods, A lot of benchmark datasets have also been introduced to standardize the model performance comparison, including NExT-QA [56], STAR [54], How2QA [31], TVQA [25], and VLEP [26]. We benchmark our network mainly on the Video Question and Answering task.
@@ 3.2 Text-guided Frame-Prompter Learning
$\pi=\left\{p_{k}\mid\frac{\exp(s_{i})}{\sum_{j=1}^{S}\exp(s_{j})},k\in[1,T/S],i,j\in[1,S]\right\},\vspace{-2mm}$ (3)
where $S$ is the frame number of one segment. Then we compute our segment mask $M_{S}$ using the segment probability $p_{k}$ and the $g_{k}$ is sampled from the $\text{Gumbel}(0,1)$ distribution,
$M_{k}=\mathcal{H}\left(\arg\max\left[g_{k}+\log p_{k}\right]\right)k\in[1,T/S],\vspace{-2mm}$ (4)
@@ 5 Conclusion, Limitation and Future Work
How do we properly ingest visual information to LLMs to utilize its capability effectively in cross-modal tasks? In this work, we present a new ViLA network with a new text-guided Frame-Prompter to smartly sample important frames, together with a cross-modal temporal distillation (QFormer-Distiller) for efficient and effective video-language alignment. From our experiments, our ViLA outperforms SOTA methods on four video question answering benchmarks and one video event prediction benchmark, especially on the temporal and interaction type of questions. We demonstrate that cross-modal keyframe selection is key to successful video-language alignment task execution. Due to resource constraints, we only evaluate on LLMs with the number of parameters not larger than 13 billions. We plan to continue research on the design of our Frame-Prompter, especially on video-language alignment for long video segments.